OpenStudy (anonymous):

how to turn 2x^2-10+8y=16x-2y^2 so it follows this form: (x-a)^2+(y-b)^2=r^2? It is for a cirkel in vectorfuctions.

OpenStudy (campbell_st):

use complete the square but 1st divide each term by 2 then rewrite the equation \[x^2 - 8x + y^2 + 8y = 5\] complete the square for the terms in x by (-8/2)^2 being added to both sides complete the square for the terms in y by adding (8/2)^2 to both sides then it will be (x +(-8/2))^2 + (y + (8/2))^2 = 5 +(-8/2)^2 + (8/2)^2 just simplify for your answer

OpenStudy (anonymous):

Close... when you divided by two... you left 8y. It should be 4y. Then subsequent fixes follow.

OpenStudy (campbell_st):

dang...simplify this (x + (-8/2))^2 + (y + (4/2)^2 = 5 + (8/2)^2 + (4/2)^2

OpenStudy (anonymous):

Perfecto!